Paul White bought 10-year, 10.8 percent coupon bonds issued by the U.S. Treasury three years ago at $902.98. If he sells these bonds, for which he paid the face value of $1,000, at the current price of $840.93, what is his realized yield on the bonds? Assume similar coupon-paying bonds make annual coupon payments. (Round intermediate calculations to 5 decimal places, e.g. 1.25145 and final answer to 2 decimal places, e.g. 15.25%.)
Answer #1
rate of return/HPR = ((Selling price+Number of Coupon amount*Coupon amount)/Purchase price-1)
=((840.93+3*108)/902.98-1)
=29.01%

realized yield = (1+HPR)^(1/n)-1

=(1+0.2901)^(1/3)-1 = 8.86%